OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (6,2) and a directix of y=1

OpenStudy (therealmeeeee):

The equation of a parabola with focus in (h,k+p) and directrix y=k-p, is, (x−h)2=4p(y−k) In the first case, h=−5k+p=−5y=k−p=7 Combining the last two relations, k=1,p=-6. So the equation of the parabola is, (x+5)2=−24(y−1)